Find the dimensions of a rectangular field that has a perimeter of 724m and an area of 32,400m^2 .
Answer by jorel555(1290)

Let w be the width of the field, and l be the length. Then:
P=2w+2l
2w+2l=724
w=362-1;
w(362-w)=32400
362w-w²=32400
w²-362w+32400=0
(w-200)(w-162)=0
w=200 or 162
The dimensions of your field are 200m by 162m. ☺☺☺☺